OpenStudy (anonymous):

Find the slope of line passing through the points (5, 9) and (5, -2) -11 0 11 undefined

OpenStudy (anonymous):

slope is (y2-y1)/(x2-x1) find it??

OpenStudy (anonymous):

undefined

OpenStudy (hoblos):

same abscissa ----> vertical line vertical lines have undifined slope

OpenStudy (anonymous):

slope = y2 - y1/(x2 - x1) = -2 - 9/(5 - 5) = -11 / 0 Therefore this ine is parallel to y axis which you can also see as x coordinate is equal. Angle = 90 degree. tan90 = Therefore undefined.